OpenStudy (gabrielah96):

put into standard form for that particular shape. 9x^2+4y^2=36

OpenStudy (mathmale):

First, do you recognize which conic section corresponds to this equation? It's not a cone, and not a parabola. What is it?

OpenStudy (gabrielah96):

How can you tell?

OpenStudy (mathmale):

We need to be able to write out and recognize the most basic equations of the four conic sections: ellipse, hyperbola, circle and parabola. In this case we know we don't have a circle because the coefficients of x^2 and y^2 are different in the expression you've typed in. In the case of a circle, the coeff. would be equal. There's more to it. But here you have an ellipse. Write out the given equation and divide every term by 36. simplify the result.

OpenStudy (gabrielah96):

so is it 1x^2/4^2+1y^2/9^2=1 ?

OpenStudy (mathmale):

\[\frac{ x^2 }{ 2^2 }+\frac{ y^2 }{ 3^2 }=1\]

OpenStudy (mathmale):

Yes, this is different from yours. Why?

OpenStudy (gabrielah96):

i didnt simplify it

OpenStudy (mathmale):

In the first term of the given equation, you have 9x^2, which you divided by 36. the result was\[\frac{ 9x^2 }{ 36 }=\frac{ x^2 }{ 4 }\] and that 4 is the same as 2^2. ok?

OpenStudy (mathmale):

Yes, and so we end up with \[\frac{ x^2 }{ 2^2 }+\frac{ y^2 }{ 3^2 }=1\] which is the equation of our ellipse in std. form.